What is the equation for the surface area of an open box? For an example a box with dimensions 2=h, 4=w, 6=l
Answer by stanbon(57953)

A box is a rectangular Prism.
Here are the instructions and an example to find the surface area.
----------------------------
Surface Area of Rectangular Prism
A rectangular prism has 2 ends and 4 sides. Opposite sides have the same area. The surface area is the sum of the areas of all six sides.
How to find the surface area of Rectangular Prisms:
Find the area of two sides (Length*Height)*2 sides
Find the area of adjacent sides (Width*Height)*2 sides
Find the area of ends (Length*Width)*2 ends
Add the three areas together to find the surface area
Example: The surface area of a rectangular prism 5 cm long, 3 cm. wide and 2 cm. high = 5*2*2 + 3*2*2 + 5*3*2 = 20 + 12 + 30 = 62 cm2.
-------------
Hope this helps.
Cheers,
Stan H.